(x^3y -^4z^-5)(x ^-4y^ -2z^9)Can some on please show me how to solve this can i use any number for the variables
thanks
angie

Answer by user_dude2008(1862)

(x^3y^-4z^-5)(x^-4y^-2z^9) = (x^3*x^-4)*(y^-4*y^-2)*(z^-5*z^9) = x^(3-4)*y^(-4-2)*z^(-5+9) = x^-1*y^-6*z^4 = z^4/(x*y^6)


Answer: (x^3y^-4z^-5)(x^-4y^-2z^9) = z^4/(x*y^6)